# Yuehao Bai

### v1

@@
A Primer on the Analysis of Randomized Experiments and a Survey of some Recent Advances thanks: The third author acknowledges support from NSF grant SES-2149408.
@@ 1 Introduction
The past two decades have witnessed a surge of new research in the analysis of randomized experiments. The emergence of this literature may seem surprising given the widespread use and long history of experiments as the “gold standard” in program evaluation, but this body of work has revealed many subtle aspects of randomized experiments that may have been previously unappreciated. This article provides an overview of some of these topics, primarily focused on stratification, regression adjustment, and cluster randomization, although we also provide short discussions on a broad range of other topics at the end of the article.
@@ 2 Setup and Notation: The Potential Outcomes Framework
We emphasize that an important feature of our discussion here and throughout the rest of the article is that we do not assume that $D^{(n)}$ are independently distributed; this will play a crucial role in our subsequent analyses when studying, for example, stratified randomization in Section 4.
@@ 3.1 Finite population analysis of $\hat{\Delta}_{n}$
This expression for the variance of $\hat{\Delta}_{n}$ is fundamental to understanding why inference on $\Delta^{\rm fp}_{N}$ is generally conservative in the finite population paradigm: note that $S^{2}_{d}$, $d\in\{0,1\}$ are simply the variances for the potential outcomes in the finite population, and we explain below how to construct estimators of these quantities. To rule out degenerate situations, we henceforth assume that at least one of $S^{2}_{d}$, $d\in\{0,1\}$ are nonzero. The quantity $S^{2}_{\Delta}$, however, is the variance of the unit-level treatment effects $(y_{j}(1)-y_{j}(0):1\leq j\leq N)$ in the finite population, and these are by definition never observed for a given unit. In settings where the experiment size $n$ is a non-trivial fraction of the population size $N$ (for instance, in design-based inference when $n=N$), this feature of the variance introduces an unavoidable roadblock for estimating $\operatorname{Var}[\hat{\Delta}_{n}]$ consistently, [^1] Here and throughout the rest of Section 3.1, consistency should be understood as saying that the ratio of the estimator to the variance converges to one in probability. and thus estimators of $\operatorname{Var}[\hat{\Delta}_{n}]$ will necessarily be conservative unless treatment effects are constant (i.e., that $y_{j}(1)-y_{j}(0)=\Delta^{\rm fp}_{N}$ for every $1\leq j\leq N$). In cases where $n$ is a vanishing fraction of $N$, however, we see that $S^{2}_{\Delta}/N$ becomes a negligible component of $\operatorname{Var}[\hat{\Delta}_{n}]$ as $N$ gets large, in the sense that in order to estimate $\operatorname{Var}[\hat{\Delta}_{n}]$ consistently it is sufficient to estimate only its first two terms consistently. As we show in Section 3.2, this feature of $\operatorname{Var}[\hat{\Delta}_{n}]$ when $n$ is a vanishing fraction of $N$ will exactly mirror our findings in the super-population analysis.
@@
$\frac{S^{2}_{1}}{n_{1}}+\frac{S^{2}_{0}}{n_{0}}-\frac{(S_{1}-S_{0})^{2}}{N}~{},$
which can be used to construct a less conservative estimator of $\operatorname{Var}[\hat{\Delta}_{n}]$. In fact, it is possible to achieve even tighter upper bounds by further exploiting the structure of $S^{2}_{\Delta}$: see Aronow et al. (2014) for details. We emphasize, however, that although we will argue in Section 3.2 that inferences based on the heteroskedasticity-robust standard error $\mathrm{SE}(\hat{\Delta}_{n})$ as in (3) will be valid under either the super-population or finite population paradigms, standard errors based on such improved upper bounds will generally be too small, and thus invalid when viewed from a super-population perspective.
@@ 3.2 Super-population analysis of $\hat{\Delta}_{n}$
It is instructive to compare the limits of the variance expressions in (2) and (5). To ensure nondegenerate limits, it is useful to scale the variances by $n$. By doing so, we see that the limit of the super-population variance mirrors the limit of the finite-population variance in a regime where we sample a vanishing fraction of the total population, i.e., $n/N\rightarrow 0$.
Since the variance of the unit-level treatment effects does not appear in our expression for $\operatorname{Var}[\hat{\Delta}_{n}]$ in (5), consistent variance estimation will be feasible in the super-population paradigm.
@@ 4 Stratified Randomized Experiments
In this section, we outline some benefits of stratification and its consequences on subsequent experimental analyses. In a stratified randomized experiment, individuals are first divided into groups (i.e., strata) sharing similar values of their baseline covariates and then assigned to treatment so as to achieve “balance” across the treatment and control groups: often, this amounts to simply performing complete randomization within each stratum. A primary motivation for stratification, going back to the work of Fisher (1935), is to ensure that the treatment and control groups are similar in the sample, in contrast to complete randomization which can only ensure that this will be true in expectation. As we will show, this property of stratification can lead to an increase in precision of the difference-in-means estimator $\hat{\Delta}_{n}$ relative to complete randomization, and as a result subsequent inferences will be unnecessarily conservative unless this is taken into consideration. In Section 4.1 we illustrate the benefits of stratification via a simple example. In Section 4.2 we discuss inference in stratified randomized experiments, including inference with “small” strata.
@@ 4.2 Inference in Stratified Experiments
Finally, we conclude with a short discussion on settings in which the assignment proportions differ across strata. This could arise, for instance, if the cost of treatment varies across different strata. We continue to assume that treatment is assigned using stratified block randomization but we let $\pi(x)$, the fraction of units that are treated in stratum $x\in\{1,2,\ldots,\mathcal{X}\}$, differ across strata. Our first observation is that the difference-in-means estimator is generally no longer consistent for $\Delta$; intuitively, the experimental design induces selection bias with respect to the stratification variable. Instead, we could consider the following estimator which computes a weighted average of the stratum-specific difference-in-means estimators, with weights determined by the strata sizes:
@@ 4.2.1 Variance Estimation with Small Strata
When $X$ was discrete, a natural estimator of $V^{\rm sbr}$ could be constructed by computing sample analogs of the means and variances of the potential outcomes at the stratum level. If, however, $X_{i}$ contains continuous components and the experiment is “finely stratified” in such a way that there is only one treated or control observation per stratum, then this logic breaks down. A leading example of when this occurs is with pair-wise matching: in a matched pairs experiment, units are paired together based on their observed covariate values and then treatment is assigned such that, in each pair, one unit is selected at random to receive treatment and the other control. With only one observation per stratum with a given treatment, we cannot estimate the stratum-level variances by simply taking sample analogs as proposed in the previous section. As discussed in Klar and Donner (1997), this added challenge to variance estimation has often been perceived as a fundamental analytical limitation of matched-pair designs. In this section, we briefly illustrate how the variance of $\hat{\Delta}_{n}$ can still be consistently estimated by using a “collapsed-strata” estimator in the spirit of Hansen et al. (1953).
@@
Intuitively, to estimate this last quantity we would want independent variation in $Y_{i}(1)$ and $Y_{i}(0)$ for each given value of $X_{i}$. Let us suppose for a moment that we did in fact have two pairs of units $\{Y_{1},Y_{2}\}$ and $\{Y_{3},Y_{4}\}$ sharing the same value of $X$. In each pair, there is exactly one treated and control observation per stratum. It follows that that $Y_{1}+Y_{2}$ and $Y_{3}+Y_{4}$ both share the same conditional mean $E[Y(1)+Y(0)|X]$. Moreover, across pairs, the outcomes are (conditionally) independent. As a consequence, we might conjecture that $(Y_{1}+Y_{2})(Y_{3}+Y_{4})$ could be used to construct an estimator for (11). Of course in practice, this scenario is not realistic, as units which are paired together will typically not share the same value of $X$. If, however, matching is performed so that we expect pairs and “adjacent” pairs to have similar characteristics, then this intuition can be formalized to construct a consistent variance estimator for (11), and thus it is possible to construct a consistent estimator of $V^{\rm sbr}$ even with “small” strata; for a formal exposition of this idea, see Bai et al. (2022, 2024b).
@@ 4.3 Further Reading
The exposition in Section 4.1 is most closely inspired by Bai (2022), who establishes the finite-sample optimality of certain matched-pair designs. Much of the discussion in Section 4.2 comes from Bugni et al. (2018), who study inference for covariate-adaptive randomization when the treated fraction is constant across strata; Bugni et al. (2019) extends the analysis to settings with multiple treatments and where the treatment proportions are allowed to differ across strata. Bai et al. (2022) studies inference for matched pair designs and develops the estimation strategy discussed in Section 4.2.1. Cytrynbaum (2023b) generalizes these estimation and inference procedures beyond pair-matching and jointly analyzes the stratification problem combined with the problem of selecting a representative sample based on covariates. Bai et al. (2024b) generalizes Bai et al. (2022) to settings with multiple treatments. Pashley and Miratrix (2021) provide an overview of design-based analyses of stratified experiments.
@@ 5.1.1 Inference on $\Delta$ when using Linear Regression Adjustment: Some Caveats
We briefly discuss some complications surrounding inference on $\Delta$ based on the regressions (12) and (15) under complete randomization, and mention some related complications when generalizing beyond complete randomization. Recall that in Section 3 we explained that the robust variance estimator obtained from the regression in (1) is consistent for the asymptotic variance of the difference-in-means estimator $\hat{\Delta}_{n}$. A natural follow-up question is then whether or not the robust variance estimators obtained from the regressions (12) and (15) are consistent for (or at least an upper bound on) the asymptotic variances of $\tilde{\beta}_{n}$ and $\hat{\beta}_{n}$, respectively. In the case of the regression in (12), the answer is yes: the robust variance estimator is consistent. In the case of the regression in (15), however, the answer is no whenever we view the sample as being drawn from a larger (finite or super-) population. The issue here is similar to our discussion surrounding consistent estimators of the variance when studying stratified randomized experiments in Section 4.2. In particular, note that the regression in (15) involves a de-meaning of the baseline covariates, and when we view the sample as being drawn from a larger population, the random fluctuations in $\bar{X}_{n}$ versus $E[X_{i}]$ need to be taken into consideration. As a result, the robust variance estimator obtained from regression (15) is not guaranteed to be valid outside of the design-based framework (in which case $\bar{X}_{n}$ is non-random). With this in mind, one simple solution discussed in Negi and Wooldridge (2021) is to instead use the built-in teffects ra command in Stata for inference.
When moving beyond complete randomization to settings with stratified assignment, the complications become more subtle. When stratification is based on a finite number of discrete categories, an easy solution is to perform regression adjustment in each stratum separately and then aggregate the stratum-level estimates; we provide relevant references for this and more complicated settings involving stratification in Section 5.3.
@@ 5.2 Double Robustness and General Regression Adjustments
where $\hat{\mu}_{1,n}(\cdot)$ and $\hat{\mu}_{0,n}(\cdot)$ are consistent estimators of the working models. $\hat{\Delta}^{\rm AIPW}_{n}$ can recover a wide range of covariate-adjusted estimators of $\Delta$ by specifying different choices of $\tilde{\mu}_{1}(\cdot)$ and $\tilde{\mu}_{0}(\cdot)$ and their corresponding estimators. For instance, if we set $\hat{\mu}_{1,n}(\cdot)=\hat{\mu}_{0,n}(\cdot)=0$ then we recover the difference-in-means estimator $\hat{\Delta}_{n}$. Alternatively, if we let $\hat{\gamma}_{n}(d)$ be the OLS estimator of the coefficient on $X_{i}$ in a regression of $Y_{i}$ on a constant and $X_{i}$ for observations with $D_{i}=d$, and set
@@
see Tu et al. (2023), Rafi (2023) for details. Importantly, this variance is the efficient variance for estimating $\Delta$ using a randomized experiment when the probability of assignment is exogenously constrained to be $\pi$ (see Armstrong, 2022; Rafi, 2023; Bai et al., 2023d).
Note that $V^{\ast}$ coincides exactly with the asymptotic variance $V^{\rm sbr}$ obtained when using the unadjusted estimator $\hat{\Delta}_{n}$ in a finely stratified randomized experiment with “small” strata (see Section 4.2.1). We have thus demonstrated two alternative methods for achieving the efficient variance $V^{*}$ when estimating $\Delta$ via a randomized experiment:
@@
This demonstrates that experiments which assign treatment using fine stratification effectively perform non-parametric regression adjustment “by design”; this feature of finely stratified experiments is further explored in Cytrynbaum (2023b) and Bai et al. (2023d).

### v2

@@
A Primer on the Analysis of Randomized Experiments and a Survey of some Recent Advances thanks: We thank the co-editor and an anonymous referee for comments and suggestions that have improved the manuscript. We would also like to thank Gustavo Bobonis for helpful discussions. The third author acknowledges support from NSF grant SES-2149408.
@@ 1 Introduction
The past two decades have witnessed a surge of new research in the analysis of randomized experiments. The emergence of this literature may seem surprising given the widespread use and long history of experiments as the “gold standard” in program evaluation, but this body of work has revealed many subtle aspects of randomized experiments that may have been previously unappreciated. This article provides an overview of some of these topics, primarily focused on stratification, regression adjustment, and cluster randomization, although we also provide short discussions on a broad range of other topics at the end of the article. We also provide a companion R package sreg, which is designed to facilitate inference in stratified (as well as potentially clustered) randomized experiments: see Trifonov et al. (2025).
@@ 2 Setup and Notation: The Potential Outcomes Framework
We emphasize that an important feature of our discussion here and throughout the rest of the article is that we do not assume that the components of $D^{(n)}$ are independently distributed; this will play a crucial role in our subsequent analyses when studying, for example, stratified randomization in Section 4.
@@ 3.1 Finite population analysis of $\hat{\Delta}_{n}$
This expression for the variance of $\hat{\Delta}_{n}$ is fundamental to understanding why inference on $\Delta^{\rm fp}_{N}$ is generally conservative in the finite population paradigm: note that $S^{2}_{d}$, $d\in\{0,1\}$ are simply the variances for the potential outcomes in the finite population, and we explain below how to construct estimators of these quantities. To rule out degenerate situations, we henceforth assume that at least one of $S^{2}_{d}$, $d\in\{0,1\}$ are nonzero. The quantity $S^{2}_{\Delta}$, however, is the variance of the unit-level treatment effects $(y_{j}(1)-y_{j}(0):1\leq j\leq N)$ in the finite population, and these are by definition never observed for a given unit. In settings where the experiment size $n$ is a non-trivial fraction of the population size $N$ (for instance, in design-based inference when $n=N$), this feature of the variance introduces an unavoidable roadblock for estimating $\operatorname{Var}[\hat{\Delta}_{n}]$ consistently, [^1] Here and throughout the rest of Section 3.1, consistency should be understood as saying that the ratio of the estimator to the variance converges to one in probability. and thus estimators of $\operatorname{Var}[\hat{\Delta}_{n}]$ will necessarily be conservative unless treatment effects are constant (i.e., that $y_{j}(1)-y_{j}(0)=\Delta^{\rm fp}_{N}$ for every $1\leq j\leq N$). In cases where $n$ is a vanishing fraction of $N$, however, we see that $S^{2}_{\Delta}/N$ becomes a negligible component of $\operatorname{Var}[\hat{\Delta}_{n}]$ as $N$ gets large, in the sense that in order to estimate $\operatorname{Var}[\hat{\Delta}_{n}]$ consistently it is sufficient to estimate only the sum of its first two terms consistently. As we show in Section 3.2, this feature of $\operatorname{Var}[\hat{\Delta}_{n}]$ when $n$ is a vanishing fraction of $N$ will exactly mirror our findings in the super-population analysis.
@@
$\frac{S^{2}_{1}}{n_{1}}+\frac{S^{2}_{0}}{n_{0}}-\frac{(S_{1}-S_{0})^{2}}{N}~{},$ (4)
which can be used to construct a less conservative estimator of $\operatorname{Var}[\hat{\Delta}_{n}]$. In fact, it is possible to achieve even tighter upper bounds by further exploiting the structure of $S^{2}_{\Delta}$: see Aronow et al. (2014) for details. We emphasize, however, that although we will argue in Section 3.2 that inferences based on the heteroskedasticity-robust standard error $\mathrm{SE}(\hat{\Delta}_{n})$ as in (3) will be valid under either the super-population or finite population paradigms, design-based standard errors (that is, standard errors which are valid in the finite population setting when $n=N$) constructed using improved upper bounds like (4) will generally be too small, and thus invalid when viewed from a super-population perspective.
@@ 3.2 Super-population analysis of $\hat{\Delta}_{n}$
It is instructive to compare the limits of the variance expressions in (2) and (6). To ensure nondegenerate limits, it is useful to scale the variances by $n$. By doing so, we see that the limit of the super-population variance mirrors the limit of the finite-population variance in a regime where we sample a vanishing fraction of the total population, i.e., $n/N\rightarrow 0$. Since the variance of the unit-level treatment effects does not appear in our expression for $\operatorname{Var}[\hat{\Delta}_{n}]$ in (6), consistent variance estimation will be feasible in the super-population paradigm.
@@ 4 Stratified Randomized Experiments
In this section, we outline some benefits of stratification and its consequences on subsequent experimental analyses. In a stratified randomized experiment, individuals are first divided into groups (i.e., strata) sharing similar values of their baseline covariates and then assigned to treatment so as to achieve “balance” across the treatment and control groups: often, this amounts to simply performing complete randomization within each stratum. Stratification is extremely common in the design of randomized experiments in all parts of the sciences (some examples in economics include Duflo et al., 2015; Berry et al., 2018; Dizon-Ross, 2019; Callen et al., 2020). A primary motivation for stratification, going back to the work of Fisher (1935), is to ensure that the treatment and control groups are similar in the sample, in contrast to complete randomization which can only ensure that this will be true in expectation. As we will show, this property of stratification can lead to an increase in precision of the difference-in-means estimator $\hat{\Delta}_{n}$ relative to complete randomization, and as a result subsequent inferences will be unnecessarily conservative unless this is taken into consideration. In Section 4.1 we illustrate the benefits of stratification via a simple example. In Section 4.2 we discuss inference in stratified randomized experiments, including inference with “small” strata.
@@ 4.2 Inference in Stratified Experiments
Finally, we conclude with a short discussion on settings in which the assignment proportions differ across strata (see, for instance, Karlan and Zinman, 2010, for an example). We continue to assume that treatment is assigned using stratified block randomization but we let $\pi(x)$, the fraction of units that are treated in stratum $x\in\{1,2,\ldots,\mathcal{X}\}$, differ across strata. Our first observation is that the difference-in-means estimator is generally no longer consistent for $\Delta$; intuitively, since the probability of assignment differs across strata, the experimental design induces selection bias with respect to the stratification variable since the composition of the covariates differs between treatment and control. Instead, we could consider the following estimator which computes a weighted average of the stratum-specific difference-in-means estimators, with weights determined by the strata sizes:
@@ 4.2.1 Variance Estimation with Small Strata
When $X$ was discrete, a natural estimator of $V^{\rm sbr}$ could be constructed by computing sample analogs of the means and variances of the potential outcomes at the stratum level. If, however, $X_{i}$ contains continuous components and the experiment is “finely stratified” in such a way that there is only one treated or control observation per stratum, then this logic breaks down. A leading example of when this occurs is with pair-wise matching: in a matched pairs experiment, units are paired together based on their observed covariate values and then treatment is assigned such that, in each pair, one unit is selected at random to receive treatment and the other control. With only one observation per stratum with a given treatment, we cannot estimate the stratum-level variances by simply taking sample analogs as proposed in the previous section. As discussed in Klar and Donner (1997), this added challenge to variance estimation has often been perceived as a fundamental analytical limitation of matched pair designs. In this section, we briefly illustrate how the variance of $\hat{\Delta}_{n}$ can still be consistently estimated by using a “collapsed-strata” estimator in the spirit of Hansen et al. (1953).
@@
Intuitively, to estimate this last quantity we would want independent variation in $Y_{i}(1)$ and $Y_{i}(0)$ for each given value of $X_{i}$. Let us suppose for a moment that we did in fact have two pairs of units $\{Y_{1},Y_{2}\}$ and $\{Y_{3},Y_{4}\}$ sharing the same value of $X$. In each pair, there is exactly one treated and control observation per stratum. It follows that that $Y_{1}+Y_{2}$ and $Y_{3}+Y_{4}$ both share the same conditional mean $E[Y(1)+Y(0)|X]$. Moreover, across pairs, the outcomes are (conditionally) independent. As a consequence, we might conjecture that $(Y_{1}+Y_{2})(Y_{3}+Y_{4})$ is approximately equal to $E[Y(1)+Y(0)|X]^{2}+(\epsilon_{1}+\epsilon_{2})(\epsilon_{3}+\epsilon_{4})$, where $E[\epsilon_{i}\epsilon_{j}|X^{(n)}]=0$ for $i$ and $j$ in different pairs. Therefore, we can construct a consistent estimator for (12) by averaging over these products across pairs of pairs. Of course in practice, this scenario is not realistic, as units which are paired together will typically not share the same value of $X$. If, however, matching is performed so that we expect pairs and “adjacent” pairs to have similar characteristics, then this intuition can be formalized to construct a consistent variance estimator for (12), and thus it is possible to construct a consistent estimator of $V^{\rm sbr}$ even with “small” strata; for a formal exposition of this idea, see Bai et al. (2022, 2024e).
@@ 4.3 Further Reading
The exposition in Section 4.1 is most closely inspired by Bai (2022), who establishes the finite-sample optimality of certain matched-pair designs. Much of the discussion in Section 4.2 comes from Bugni et al. (2018), who study inference for covariate-adaptive randomization when the treated fraction is constant across strata; Bugni et al. (2019) extends the analysis to settings with multiple treatments and where the treatment proportions are allowed to differ across strata. Bai et al. (2022) studies inference for matched pair designs and develops the estimation strategy discussed in Section 4.2.1. Cytrynbaum (2023b) generalizes these estimation and inference procedures beyond pair-matching and jointly analyzes the stratification problem combined with the problem of selecting a representative sample based on covariates. Bai et al. (2024e) generalizes Bai et al. (2022) to settings with multiple treatments. Pashley and Miratrix (2021) and Bai et al. (2025a) provide overviews of design-based analyses of stratified experiments.
@@ 5.1.1 Inference on $\Delta$ when using Linear Regression Adjustment: Some Caveats
We briefly discuss some complications surrounding inference on $\Delta$ based on the regressions (13) and (16) under complete randomization, and mention some related complications when generalizing beyond complete randomization. Recall that in Section 3 we explained that the robust variance estimator obtained from the regression in (1) is consistent for the asymptotic variance of the difference-in-means estimator $\hat{\Delta}_{n}$. A natural follow-up question is then whether or not the robust variance estimators obtained from the regressions (13) and (16) are consistent for (or at least an upper bound on) the asymptotic variances of $\tilde{\beta}_{n}$ and $\hat{\beta}_{n}$, respectively. In the case of the regression in (13), the answer is yes: the robust variance estimator is consistent because $\tilde{\beta}_{n}$ is simply the OLS estimator of the best linear predictor of $Y$ given $1,D,$ and $X$. In the case of the regression in (16), however, the answer is no whenever we view the sample as being drawn from a larger (finite or super-) population. The issue here is similar to our discussion surrounding consistent estimators of the variance when studying stratified randomized experiments in Section 4.2. In particular, note that the regression in (16) involves a de-meaning of the baseline covariates, and when we view the sample as being drawn from a larger population, the random fluctuations in $\bar{X}_{n}$ versus $E[X_{i}]$ need to be taken into consideration. As a result, the robust variance estimator obtained from regression (16) is not guaranteed to be valid outside of the design-based framework (in which case $\bar{X}_{n}$ is non-random). When moving beyond complete randomization to settings with stratified assignment, the complications become more subtle. When stratification is based on a finite number of discrete categories, an easy solution is to perform regression adjustment in each stratum separately and then aggregate the stratum-level estimates; we provide relevant references for this and more complicated settings involving stratification in Section 5.3. With these considerations in mind, we suggest that practitioners use our R command sreg to perform regression adjustment in (stratified) randomized experiments: see Trifonov et al. (2025).
@@ 5.2 Double Robustness and General Regression Adjustments
where $\hat{\mu}_{1,n}(\cdot)$ and $\hat{\mu}_{0,n}(\cdot)$ are consistent estimators of the working models. $\hat{\Delta}^{\rm AIPW}_{n}$ can recover a wide range of covariate-adjusted estimators of $\Delta$ by specifying different choices of $\tilde{\mu}_{1}(\cdot)$ and $\tilde{\mu}_{0}(\cdot)$ and their corresponding estimators. For instance, if we set $\hat{\mu}_{d,n}=\frac{1}{n_{d}}\sum_{1\leq i\leq n}Y_{i}I\{D_{i}=d\}$, then we recover the difference-in-means estimator $\hat{\Delta}_{n}$ (if we instead set $\hat{\mu}_{1,n}(\cdot)=\hat{\mu}_{0,n}(\cdot)=0$, then we recover the Horvitz-Thompson estimator of the average treatment effect). Alternatively, if we let $\hat{\gamma}_{n}(d)$ be the OLS estimator of the coefficient on $X_{i}$ in a regression of $Y_{i}$ on a constant and $X_{i}$ for observations with $D_{i}=d$, and set
@@
see Tu et al. (2023), Rafi (2023) for details. Importantly, this variance is the efficient variance for estimating $\Delta$ using a randomized experiment when the probability of assignment is exogenously constrained to be $\pi$ (see Armstrong, 2022; Rafi, 2023; Bai et al., 2025b).
Note that $V^{\ast}$ coincides exactly with the asymptotic variance $V^{\rm sbr}$ obtained when using the unadjusted estimator $\hat{\Delta}_{n}$ in a finely stratified randomized experiment with “small” strata (see Section 4.2.1). We have thus demonstrated two alternative methods for achieving the efficient variance $V^{*}$ when estimating $\Delta$ via a randomized experiment: [^3] Of course, this list of methods is not exhaustive. For instance, efficient estimation could also be achieved under i.i.d. randomization using a standard inverse-propensity weighted (IPW) estimator, where the propensity score is non-parametrically estimated as a function of $X$: see in particular Hirano et al. (2003) for details.
@@
This demonstrates that experiments which assign treatment using fine stratification effectively perform non-parametric regression adjustment “by design”; this feature of finely stratified experiments is further explored in Cytrynbaum (2023b) and Bai et al. (2025b). In principle, both strategies could be combined, by first stratifying on the covariates which might be expected to be most predictive for the outcome of interest, and then adjusting for a (potentially high-dimensional) set of additional covariates via regression ex-post: see Bai et al. (2024c) for details.